OpenStudy (anonymous):

Kevins age can be represented by the expression 12p^3. Jakes age can be represented by the expression 4p^5 What is the ratio of Kevins age to Jakes age?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@AmberLOVESyou!.

OpenStudy (igreen):

12p^3/4p^5

OpenStudy (igreen):

Simplify Divide the whole numbers and subtract the exponents

OpenStudy (anonymous):

3/ 4^2

OpenStudy (anonymous):

3/ 4p^2

OpenStudy (igreen):

No.. 12 / 4 = ? 3 - 5 = ?

OpenStudy (anonymous):

-2

OpenStudy (igreen):

12 / 4 = ?